OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x+c)^2=a. 10x^2+80x+220=0 Can someone please explain this?

OpenStudy (raden):

It's pretty, but hard to teach first, u need simplify of 10x^2+80x+220=0 divided by 10 to both sides, giving us x^2+8x+22=0 now, to get a perpect square form, setting become x^2 + 8x + 16 + 6 = 0 or x^2 + 8x + 16 = -6 (x+4)^2 = -6 sorry, if my explain not make sense

OpenStudy (anonymous):

No. Most of it makes sense. But I don't get how you find the number for c.

OpenStudy (raden):

if given (x+4)^2 = -6 compare the general form above : (x+c)^2=a it mean, c=4 and a=-6